OpenStudy (anonymous):

help please? Part 1: Explain which companies would advertise APY over APR and describe how this would help them attract customers. Part 2: Create a unique APR (state how often the rate is compounded) and calculate the corresponding APY. Use a comparison of the two rates to verify your answer to part 1.

OpenStudy (amistre64):

your lacking information for the problem in your post

OpenStudy (amistre64):

APR of $1, compounded is:\[(1+\frac rn)^n\] APY is the rate it would have been if it had been stated at a compounding of 1 year:\[(1+k)\]where k is the APY \[1+k=(1+\frac rn)^n\] \[k=(1+\frac rn)^n-1\]

OpenStudy (anonymous):

Yes it is a bit vague isn't it? Would this be the answer to my question or a formula to plug my own numbers into?

OpenStudy (amistre64):

this is a general formula to aid you along the way let r be the rate that is compounded, let n be the number of times a year it gets assessed, and then k (or the APY) is just a function r and n as stated

OpenStudy (amistre64):

since the APR and the APY give the same "values" at the end of the year, the strategy lies in how people relate the value of a rate to their investments

OpenStudy (anonymous):

I am supposed to create my own scenaario here. But I can't come up with a rational one to really protray the relationship between apr and apy.

OpenStudy (amistre64):

usually, a higher numerical value attracts more investors on a psychological level

OpenStudy (amistre64):

think of a rate ... think of how many times a year you want to assess that rate ...

OpenStudy (anonymous):

okay so. 15.5% compounded monthly

OpenStudy (amistre64):

then r = 15.5, n=12 and k = (1+.0155/12)^12 - 1

OpenStudy (amistre64):

got my decie in the wrong spot :) 0.155/12 that is

OpenStudy (anonymous):

.0129

OpenStudy (amistre64):

APR = 15.5 % APY = 16.6449 % according to the workings of my patent pending formulas

OpenStudy (anonymous):

Ah okay, this seems to make it more concrete. thank you :)

OpenStudy (amistre64):

youre welcome; and just to fix a typo ... APY = 16.6499 % these old eyes aint what they used ta be

OpenStudy (anonymous):

Ahaha, I was checking that on my calculator and I noticed that too, no worries. It happens to the best of us!